# RGB-Only Supervised Camera Parameter Optimization in Dynamic Scenes

### v1

@@ Abstract
Although COLMAP has long remained the predominant method for camera parameter optimization in static scenes, it is constrained by its lengthy runtime and reliance on ground truth (GT) motion masks for application to dynamic scenes. Many efforts attempted to improve it by incorporating more priors as supervision such as GT focal length, motion masks, 3D point clouds, camera poses, and metric depth, which, however, are typically unavailable in casually captured RGB videos. In this paper, we propose a novel method for more accurate and efficient camera parameter optimization in dynamic scenes solely supervised by a single RGB video. Our method consists of three key components: (1) Patch-wise Tracking Filters, to establish robust and maximally sparse hinge-like relations across the RGB video. (2) Outlier-aware Joint Optimization, for efficient camera parameter optimization by adaptive down-weighting of moving outliers, without reliance on motion priors. (3) A Two-stage Optimization Strategy, to enhance stability and optimization speed by a trade-off between the Softplus limits and convex minima in losses. We visually and numerically evaluate our camera estimates. To further validate accuracy, we feed the camera estimates into a 4D reconstruction method and assess the resulting 3D scenes, and rendered 2D RGB and depth maps. We perform experiments on 4 real-world datasets (NeRF-DS, DAVIS, iPhone, and TUM-dynamics) and 1 synthetic dataset (MPI-Sintel), demonstrating that our method estimates camera parameters more efficiently and accurately with a single RGB video as the only supervision.
@@ 1 Introduction
Based on these insights, we propose an RGB-only supervised, accurate, and efficient camera parameter optimization method, with a brief performance overview in Figure˜1. Specifically, to minimize reliance on pre-trained dense prediction models while still establishing robust and maximally sparse hinge-like relations across the video as accurate pseudo-supervision (bottom right corner in Figure˜2), we propose the novel patch-wise tracking filters built solely on a pre-trained point tracking (PT) model. This formulation effectively avoids inaccurate tracking trajectories extracted across frames and computational latency induced by the noisy dense prediction as pseudo-supervision.
@@ 3.1 Patch-wise Tracking Filters
Built on a pre-trained PT model, we observe that its attention mechanism assigns higher attention weights to pixels with more accurate tracking results which are always texture-rich pixels with large gradient norms. Inspired by it, as shown in Figure˜2, we propose the patch-wise texture filter to identify the high-texture patches within $F_{t}$ and the patch-wise gradient filter to select the pixel with the highest gradient norm within each identified patch. While tracking such identified points, the visibility filter keeps removing trajectories that become invisible and the patch-wise distribution filter keeps the one with the largest gradient norm when multiple moving points enter the same patch. As shown in section˜E.2.1 (fig.˜9), our method only retains the robust and accurate trajectories as pseudo-supervision.
@@ 4.2 Time Efficiency Evaluation
Figure 5: Qualitative NVS Results on DAVIS davis . Our performance is the best because of our accurate camera estimates. More are in section˜E.2.2 (fig.˜13, fig.˜14, fig.˜15, and fig.˜16).
@@ 4.4 NVS Evaluation
Figure 7: Qualitative NVS Results on iPhone iphone . Our method outperforms other SOTA RGB-only supervised approaches and even surpasses LiDAR-based Record3D when the movement in scenes with large motion (top row). More are in section˜E.2.2 (fig.˜10, fig.˜11, and fig.˜12).
@@
Figure 8: Qualitative NVS results on NeRF-DS nerfds . Our renderings are the most plausible. More are in section˜E.2.2 (fig.˜17).
@@ 5 Conclusion and Limitation
We proposed a new RGB-only supervised, accurate, and efficient camera parameter optimization method in casually collected dynamic-scene videos. Our method effectively tackles the challenge of precisely and efficiently estimating per-frame camera parameters under the situation of having no additional GT supervision (e.g. GT motion masks, focal length, 3D point clouds, metric depth, and camera poses) other than RGB videos, which is the most common scenario in real-world and consumer-grade reconstructions. Our method may serve as a step towards high-fidelity dynamic scene reconstruction from casually captured videos. Due to the space limit, the limitation is discussed in appendix˜F.
@@ E.2.1 Trajectories from Patch-wise Tracking Filters as Pseudo-supervision
In Figure˜9, we show the trajectory comparisons on the NeRF-DS nerfds dataset as samples. As discussed in the 3rd paragraph in section˜1, and section˜3.1 of the main paper, our patch-wise tracking filters can establish robust and maximally sparse hinge-like relations as accurate pseudo-supervision, avoiding noisy and inaccurate tracking trajectories. In Figure˜9, it is easy to see our proposed method avoids the inaccurate ones in the low-texture regions (walls), and meanwhile, adaptively adds new reliable trajectories when the number of left trajectories on each frame is less than $B$.
@@ E.2.2 NVS
We show more RGB and depth rendering results on NeRF-DS nerfds , DAVIS davis , and iPhone iphone dataset in Figure˜10, Figure˜11, Figure˜12, Figure˜13, Figure˜14, Figure˜15, Figure˜16, and Figure˜17. It is easy to see that the RGB and depth rendering results of our method are better than other RGB-only supervised approaches. In addition, the performance of our method is also comparable with that of the LiDAR-based Record3D app.
@@ E.2.3 Optimized 3D Gaussian Fields
Since the iPhone iphone dataset is the most challenging dataset with large camera and object movements, we show more visualizations of optimized 3D Gaussian fields in Figure˜19, Figure˜20, and Figure˜18. Such comparisons demonstrate that our camera estimates enable superior reconstruction of 3D Gaussian fields compared to other RGB-only supervised approaches. Moreover, the reconstructed fields using our estimates are comparable to, or even surpass, those obtained with the LiDAR-based Record3D record3d app, particularly in scenes with significant motion.
@@
Appendix F Limitation
Although our proposed method is currently the most state-of-the-art RGB-only supervised, accurate, and efficient camera parameter optimization method in dynamic scenes, there are still several limitations. We assume a constant focal length throughout the video. While this assumption is reasonable and currently common to SOTA, the task of accurate and efficient camera parameter optimization for dynamic scene videos with zooming effects under RGB-only supervision remains an open problem. Another common challenge for RGB-only supervised methods, not addressed in this paper, is maintaining robustness in scenes dominated by large moving objects. As shown in fig.˜21, the screen space is occupied by the moving human and dragon. It is challenging for our method to establish robust and maximally sparse hinge-like relations as accurate pseudo-supervision because most of the extracted trajectories belong to outliers. CasualSAM casualsam struggles due to the rapid changes in depth maps from frame to frame, making 3D space alignment difficult. We plan to maintain consistency in our input setup and address these challenges as part of future research.

### v2

@@ Abstract
Although COLMAP has long remained the predominant method for camera parameter optimization in static scenes, it is constrained by its lengthy runtime and reliance on ground truth (GT) motion masks for application to dynamic scenes. Many efforts attempted to improve it by incorporating more priors as supervision such as GT focal length, motion masks, 3D point clouds, camera poses, and metric depth, which, however, are typically unavailable in casually captured RGB videos. In this paper, we propose a novel method for more accurate and efficient camera parameter optimization in dynamic scenes solely supervised by a single RGB video, dubbed ROS-Cam. Our method consists of three key components: (1) Patch-wise Tracking Filters, to establish robust and maximally sparse hinge-like relations across the RGB video. (2) Outlier-aware Joint Optimization, for efficient camera parameter optimization by adaptive down-weighting of moving outliers, without reliance on motion priors. (3) A Two-stage Optimization Strategy, to enhance stability and optimization speed by a trade-off between the Softplus limits and convex minima in losses. We visually and numerically evaluate our camera estimates. To further validate accuracy, we feed the camera estimates into a 4D reconstruction method and assess the resulting 3D scenes, and rendered 2D RGB and depth maps. We perform experiments on 4 real-world datasets (NeRF-DS, DAVIS, iPhone, and TUM-dynamics) and 1 synthetic dataset (MPI-Sintel), demonstrating that our method estimates camera parameters more efficiently and accurately with a single RGB video as the only supervision.
@@ 1 Introduction
Based on these insights, we propose ROS-Cam, an RGB-only supervised, accurate, and efficient camera parameter optimization method, with a brief performance overview in Figure˜1. Specifically, to minimize reliance on pre-trained dense prediction models while still establishing robust and maximally sparse hinge-like relations across the video as accurate pseudo-supervision (bottom right corner in Figure˜2), we propose the novel patch-wise tracking filters built solely on a pre-trained point tracking (PT) model. This formulation effectively avoids inaccurate tracking trajectories extracted across frames and computational latency induced by the noisy dense prediction as pseudo-supervision.
@@ 3.1 Patch-wise Tracking Filters
Built on a pre-trained PT model, we observe that its attention mechanism assigns higher attention weights to pixels with more accurate tracking results which are always texture-rich pixels with large gradient norms. Inspired by it, as shown in Figure˜2, we propose the patch-wise texture filter to identify the high-texture patches within $F_{t}$ and the patch-wise gradient filter to select the pixel with the highest gradient norm within each identified patch. While tracking such identified points, the visibility filter keeps removing trajectories that become invisible and the patch-wise distribution filter keeps the one with the largest gradient norm when multiple moving points enter the same patch. As shown in section˜E.2.1 (fig.˜10), our method only retains the robust and accurate trajectories as pseudo-supervision.
@@ 4.2 Time Efficiency Evaluation
Figure 5: Qualitative NVS Results on DAVIS davis . Our performance is the best because of our accurate camera estimates. More are in section˜E.2.2 (fig.˜14, fig.˜15, fig.˜16, and fig.˜17).
@@ 4.4 NVS Evaluation
Figure 7: Qualitative NVS Results on iPhone iphone . Our method outperforms other SOTA RGB-only supervised approaches and even surpasses LiDAR-based Record3D when the movement in scenes with large motion (top row). More are in section˜E.2.2 (fig.˜11, fig.˜12, and fig.˜13).
@@
Figure 8: Qualitative NVS results on NeRF-DS nerfds . Our renderings are the most plausible. More are in section˜E.2.2 (fig.˜18).
@@ 5 Conclusion and Limitation
We proposed a new RGB-only supervised, accurate, and efficient camera parameter optimization method in casually collected dynamic-scene videos. Our method effectively tackles the challenge of precisely and efficiently estimating per-frame camera parameters under the situation of having no additional GT supervision (e.g. GT motion masks, focal length, 3D point clouds, metric depth, and camera poses) other than RGB videos, which is the most common scenario in real-world and consumer-grade reconstructions. Our method may serve as a step towards high-fidelity dynamic scene reconstruction from casually captured videos.
Although our proposed method is currently the most state-of-the-art RGB-only supervised, accurate, and efficient camera parameter optimization method in dynamic scenes, there are still several limitations. We assume a constant focal length throughout the video. While this assumption is reasonable and currently common to SOTA, the task of accurate and efficient camera parameter optimization for dynamic scene videos with zooming effects under RGB-only supervision remains an open problem. Another common challenge for RGB-only supervised methods, not addressed in this paper, is maintaining robustness in scenes dominated by large moving objects. As shown in fig.˜9, the screen space is occupied by the moving human and dragon. It is challenging for our method to establish robust and maximally sparse hinge-like relations as accurate pseudo-supervision because most of the extracted trajectories belong to outliers. CasualSAM casualsam struggles due to the rapid changes in depth maps from frame to frame, making 3D space alignment difficult. We plan to maintain consistency in our input setup and address these challenges as part of future research.
@@ E.2.1 Trajectories from Patch-wise Tracking Filters as Pseudo-supervision
In Figure˜10, we show the trajectory comparisons on the NeRF-DS nerfds dataset as samples. As discussed in the 3rd paragraph in section˜1, and section˜3.1 of the main paper, our patch-wise tracking filters can establish robust and maximally sparse hinge-like relations as accurate pseudo-supervision, avoiding noisy and inaccurate tracking trajectories. In Figure˜10, it is easy to see our proposed method avoids the inaccurate ones in the low-texture regions (walls), and meanwhile, adaptively adds new reliable trajectories when the number of left trajectories on each frame is less than $B$.
@@ E.2.2 NVS
We show more RGB and depth rendering results on NeRF-DS nerfds , DAVIS davis , and iPhone iphone dataset in Figure˜11, Figure˜12, Figure˜13, Figure˜14, Figure˜15, Figure˜16, Figure˜17, and Figure˜18. It is easy to see that the RGB and depth rendering results of our method are better than other RGB-only supervised approaches. In addition, the performance of our method is also comparable with that of the LiDAR-based Record3D app.
@@ E.2.3 Optimized 3D Gaussian Fields
Since the iPhone iphone dataset is the most challenging dataset with large camera and object movements, we show more visualizations of optimized 3D Gaussian fields in Figure˜20, Figure˜21, and Figure˜19. Such comparisons demonstrate that our camera estimates enable superior reconstruction of 3D Gaussian fields compared to other RGB-only supervised approaches. Moreover, the reconstructed fields using our estimates are comparable to, or even surpass, those obtained with the LiDAR-based Record3D record3d app, particularly in scenes with significant motion.